Which of the following species are/is present at the highest concentration in a 1.0 mol/L aqueous HCN solution?

a) CN-   b) H3O+   c) HCN d) OH-   e) a) and b) are present at the same concentration, which is greater than that of the other listed species.
Answer #1

        HCN(aq) +H2O----> H3O^+ (aq) + CN^- (aq)

I       1                          0                     0

C      -x                             +x                 +x

E     1-x                          +x                   +x

            ka    = [H^+][CN^-]/[HCN]

             6.2*10^-10 = x*x/(1-x)

              6.2*10^-10*(1-x) = x^2

                x = 2.5*10^-5

      [H3O^+]   = x   = 2.5*10^-5 M

      [CN^-]   = x   = 2.5*10^-5 M

     [HCN]   = 1-x   = 1-2.5*10^-5   = 0.999975M has highest concentration

    [OH^-]    = Kw/[H3O^+]

                   = 1*10^-14/(2.5*10^-5)   = 4*10^-10 M

c) HCN >>>>answer